height of hcp unit cell


 

Grade:12

1 Answers

Aman Bansal
12 years ago

Dear Mukal,

A,B,C,E (red dots) will represents the centres of sphere of atoms in HCP closed packing,

now, AB=BC=CA=EB = 2R (edge lengths of HCP)

BD=\frac{2R}{\sqrt{3}}  ,as BD divides the median through B in 2:1 ratio from B...

now EDB is rt. triangle,

by pythagores theorem :

(\frac{2R}{\sqrt{3}})^2+h^2=(2R)^2

=> h=2R\sqrt{\frac{2}{3}}

now, their will be another atom which will be mirror image of E on plane ABC

so, the height (c) of HCP will be 2h

c=2h=4R\sqrt{\frac{2}{3}}

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert